# Conventional and practical metallic superconductivity arising from repulsive Coulomb coupling

### v1

@@ Abstract
A concrete question is discussed: Can there be conventional $s$-wave superconductivity in regular 3D metals, i.e., electrons in a jellium background, interacting via the standard Coulomb coupling? We are interested in ‘practical’ superconductivity that can in principle be observed in experiments, so the $T=0$ ground state being superconducting is not of interest, or for that matter a $T_{c}$ which is exponentially small and therefore ‘impractical’ is also not of interest in the current work. We find that almost any theory based on the BCS-Migdal-Eliashberg paradigm, with some form of screened Coulomb coupling replacing the electron-phonon coupling in the BCS or Eliashberg theory, would uncritically predict absurdly high $T_{c}\sim 100$ K in all metals (including the alkali metals, which are well-described by the jellium model) arising from the unavoidable fact that the Fermi, plasmon, and Coulomb potential energy scales are all $10^{4}$ K. Therefore, we conclude, based on reduction ad absurdum, that the violation of the venerable Migdal theorem in this problem is sufficiently disruptive that no significance can be attached to numerous existing theoretical publications in the literature claiming plasmon-induced (or other similar Coulomb coupling-induced) practical SC. Using a careful analysis of the Eliashberg gap equations we find that the superconducting $T_{c}$ of the 3D electron gas can be reduced below the $\sim 1$ K range depending on choices of frequency and momentum cut-off parameters that are introduced to satisfy Migdall’s theorem but are apriori unknown. The only believable result is the one discovered sixty years ago [1] by Kohn and Luttinger predicting non-$s$-wave SC arising from Friedel oscillations with exponentially (and unobservably) low $T_{c}$. We provide several theoretical approaches using both BCS and Eliashberg theories and different screening models to make our point.
@@ I Introduction
Superconductivity is a ubiquitous phenomenon which has remained central to physics for more than 100 years ever since its discovery in 1911 [2]. There are currently roughly $\sim 15{,}000$ known superconductors with $T_{c}$ ranging from $\sim 10$ mK to $\sim 250$ K (under high pressure). The search for new superconducting materials with practical use is an active research area in materials science with many promises for quantum computers and large language model based artificial intelligence revolutionizing such efforts often hyped in the media. On a fundamental level, superconductivity, being a quintessential example of spontaneous symmetry breaking (namely, the $U(1)$ symmetry), continues to attract great attention from theorists and experimentalists alike with the ongoing recent focus being room temperature superconductors, topological superconductors, and exotic superconductors with unusual order parameter symmetries. It may not be an exaggeration to say that superconductivity may very well be the most actively studied single topic in all of physics over the last 100 or so years.
@@
In this context, one issue that has fascinated physicists a great deal over the years is the superconducting mechanism—what leads to the pairing glue causing the superconducting instability? The minimal physical picture for superconductivity, and this is the relevant physics for the current work too, is that electrons pair up (“Cooper pairs” [3]) because of an attractive interaction coupling them at the Fermi level causing the exchange of a virtual boson which acts as the glue to create the Cooper pairs. The glued Cooper pairs then condense into a zero resistance collective ground state producing the breaking of $U(1)$ symmetry since the Cooper pairs form the condensate effectively breaking the gauge invariance. The key physics is, however, the pairing arising from an effective attractive interaction between opposite spin electrons at the Fermi surface. Thus, a mechanism is needed to cause an effective attractive interaction between electrons near the Fermi surface. For most superconductors (if not all) this pairing arises from the electron-phonon interaction, which leads to a phonon-mediated effective attraction between the electron pairs with opposite spins near the Fermi surface. Since the electron spins are opposite, the spin part of the paired wavefunction is antisymmetric, making the spatial symmetry to be symmetric, and hence a singlet $s$-wave superconductivity. All superconducting metals are such spin-singlet spatially symmetric $s$-wave superconductors, with phonons being the bosonic glue between the electron pairs. In fact, one could probably claim that there is no known superconductor which has been compellingly shown to have a pairing glue other than electron-phonon coupling although there certainly are candidate materials where the situation remains an open question such as high-$T_{c}$ cuprate superconductors, heavy fermion superconductors, iron based pnictide superconductors, and perhaps various superconductors arising in multilayer graphene-type systems.
@@
The purpose of the current work, which is partly a perspective, is to investigate critically the extent to which the electron-electron interaction by itself could lead to any ‘practical’ $s$-wave superconductivity without the presence of any electron-phonon interaction (or other bosonic coupling outside the scope of just the electron system itself) in the Hamiltonian. At first, the question seems absurd since the basic electron-electron interaction, i.e., Coulomb coupling, is by definition repulsive, seemingly ruling out prima facie any attractive glue for pairing, let alone the formation of a superconducting condensate of the pairs. It turns out, however, that the situation is subtle since electrons screen each other, and the actual interaction between electrons is not the bare Coulomb interaction, but the many-body screened interaction, whose structure in general is extremely complicated. A seminal paper [1] by Kohn and Luttinger (KL SC) pointed out a long time ago that the ground state of an interacting electron liquid, at least for a weak short-range screened interaction, is superconducting, hinting at the possibility that interacting electron liquids may very well manifest superconductivity arising just from the Coulomb coupling. But the KL SC is neither the conventional $s$-wave SC nor ‘practical’ by any stretch of imagination, as the estimated $T_{c}$ is exponentially low for real metals ($T_{c}\sim 10^{-10}$ to $10^{-100}$ K), and occurs only for higher orbital angular momenta. Thus, KL SC is only of academic interest, establishing a matter of principle that indeed repulsive Coulomb coupling under some circumstances can lead to a superconducting ground state, albeit a very fragile unconventional one of no practical interest. KL SC arises from the Friedel oscillations in screening associated with the existence of the Fermi surface which leads to $2k_{F}$-periodic oscillatory behavior in the statically screened Coulomb interaction, and as such can only occur not for the spherically symmetric $s$-wave SC, but for higher angular momentum orbital symmetry such $p$, $d$ wave SC. There has been renewed interest in KL SC in the context of lattice systems such as the Hubbard model [4] and 2D moiré materials [5] where the existence of van Hove singularities may enhance the KL SC $T_{c}$, but it is still low and the SC is always in higher orbital angular momentum channel.
@@
So, the KL SC, while establishing a matter of principle in the existence of SC via a strictly non-phonon mechanism, does not address the question we pose in this work: Can there be conventional $s$-wave metallic superconductivity which is also practical (i.e. experimentally observable with a reasonable $T_{c}$)? Much work has been done since the 1980s in the possibility of the so-called ‘plasmon-mediated’ metallic superconductivity in electron liquids, with claims of high $T_{c}$ $s$-wave metallic superconductivity arising from only electron-electron interactions. The basic physical idea is deceptively simple and appealing: Plasmons are quantized bosonic collective modes of the well-known ‘plasma wave’ charge density oscillations arising from the long-range Coulomb coupling, and the electron-plasmon interaction may act similarly to the electron-phonon interaction providing the pairing glue, leading to metallic superconductivity. If so, the putative $T_{c}$ of such plasmon-induced superconductivity could be very high since the energy scales of all electronic excitations in metals are extremely high, $10^{4}$ K in contrast to phonons whose characteristic energy scale is $\sim 10^{2}$ K. So, by just changing the bosonic glue from phonons to plasmons, it seems plausible that the $T_{c}$ could be raised by >2>2 orders of magnitude! In fact, the cuprate SCs with $T_{c}\sim 100$ K are thus heuristically consistent with an effective electron-electron induced SC since their $T_{c}$ is crudely a factor of 100 larger than the corresponding $T_{c}\sim 1$$10$ K in conventional metals such as Al and Pb which are known to manifest phonon-mediated SC for sure. This in fact has motivated people to propose the plasmon mechanism for the cuprate superconductivity although this is definitely a rather small minority view by no means accepted in the larger high-$T_{c}$ community [6, 7, 8, 9]. The crucial problem in such suggestions for plasmon mediated superconductivity in electron liquids is the stark absence of generic high-temperature ($T_{c}>100$ K) metallic superconductivity in normal metals which all have plasmons with energy scales $10^{4}$ K and Fermi energies also $10^{4}$ K. If plasmons and electron-electron interactions could generically lead to metallic superconductivity, Na and K and Li should all be room temperature superconductors, but factually, they are not superconductors at all under ambient conditions in spite of being ideal jellium electron liquids with electronic energy scales $\sim 5\times 10^{4}$ K. But, all known (>50>50 different elements, some under high pressure) metallic superconductors with plasmon energies $10^{4}$ K, are known to be phonon-mediated $s$-wave SCs with $T_{c}$ (mostly) $\sim 1$$10$ K which is in line with the typical phononic energy scale $\sim 10^{2}$ K. In fact, a major problem with the plasmon-mediated (or other similar) electron interaction mediated SC proposal is: Why is the $T_{c}$ so low in all such SCs (even in high-$T_{c}$ cuprates, $T_{c}\sim 100$ K, orders of magnitude below the electronic kinetic energy and interaction energy scales)?
@@
We show in this work, using plausible BCS theory based arguments, that the generically expected BCS $T_{c}$ for the plasmon mediated metallic SC is $\sim 100$$1000$ K, an absurdly high number when talking about conventional metallic superconductors which have been experimentally studied extensively for almost 120 years. Obviously, this ‘plausible’ straightforward BCS theory, with the plasmon modes replacing the phonon modes and the electron-plasmon coupling replacing the electron-phonon coupling, does not work as it predicts manifestly incorrect $T_{c}$ which is incompatible with reality. Alkali metals are not $\sim 100$ K superconductors, and the superconducting metals get their superconductivity from electron-phonon interactions, not the electron-electron interactions. Therefore, the question arises what is wrong with the BCS theory applied to plasmons compared with that for phonons where generally it has had predictive success in producing $T_{c}$ and superconducting gaps in agreement with experiments in numerous situations [10, 11]. We discuss this point from several critical perspectives using the Eliashberg theory, emphasizing that it is possible to predict essentially any $T_{c}$ based on the BCS-Eliashberg theory for plasmon-induced superconductivity since the theory is uncontrolled because of the crucial inapplicability of Migdal theorem for the electron-plasmon interaction [12]. The successful BCS-Eliashberg theory for superconductivity, enabling a reasonable quantitative $T_{c}$ prediction for metallic superconductors, is grounded on the Migdal theorem which applies to the electron-phonon interaction problem in metals by virtue of the Fermi energy being much larger than the typical phonon energy (the so-called Debye temperature), but not for the electron-electron interaction induced plasmonic superconductivity.
@@
Now we immediately face a serious conundrum in discussing superconductivity induced by plasmons. Since the plasmons arise intrinsically as collective modes of the interacting electron system, unlike the phonons which arise from the lattice vibrations, there cannot be, by definition, any Migdal theorem for the electron-plasmon interaction. Indeed, the plasmon energy in metals ($\sim 10$ eV) is comparable to the Fermi energy $E_{F}$, and is also similar to the typical average metallic inter-electron Coulomb potential energy. Thus, there is no energy scale separation (or velocity scale separation) enabling the neglect of vertex corrections. Hence, the use of BCS-Eliashberg theory for the calculation of $T_{c}$ in the plasmon induced metallic superconductivity is fundamentally flawed as one must include vertex corrections to all orders in the theory, which is obviously impossible. We show that it is easy to obtain $T_{c}\sim 100$$1000$ K in metals arising from electron-plasmon interaction simply by following the standard BCS-Eliashberg-Migdal theory (with plasmons replacing phonons), but these results are unreliable since the neglect of vertex corrections in the theory is unjustified and uncontrolled. Similarly, the estimate of $\mu^{*}$ in the plasmon case is nontrivial and ambiguous since one is starting entirely with a repulsive interaction. We elaborate on these subtle but important theoretical issues in the rest of this paper.
The rest of this paper is organized as follows. In Sec. II, we provide the calculated dynamically screened Coulomb interaction in 3D jellium metals for several many-body approximations to show explicitly that indeed the effective screened interaction is attractive in large regimes of the energy-momentum space, making a discussion of possible superconductivity arising strictly from electron-electron interactions meaningful. In Sec. III, we use the effective electron-plasmon interaction from Sec. II to obtain the SC $T_{c}$ induced by plasmons in the standard BCS theory, finding $T_{c}\sim 100$ K in typical metals (which is in stark disagreement with experiments). In Sec. IV, we develop a full Eliashberg-Migdal theory, and point out precisely why the theory is unreliable in predicting $T_{c}$ by discussing the subtle frequency-dependent retardation effects which are crucial in the Eliashberg theory. We conclude in Sec. V, summarizing our findings and emphasizing that no compelling theoretical case can be made that electron-plasmon interaction can or should lead to conventional and practical superconductivity in normal metals, although uncontrolled approximations do typically lead to $T_{c}\sim 100$ K for plasmon-induced metallic superconductivity.

### v2

@@ Abstract
A concrete question is discussed: Can there be conventional $s$-wave superconductivity in regular 3D (or 2D) metals, i.e., electrons in a jellium background, interacting via the standard Coulomb coupling? We are interested in ‘practical’ superconductivity that can in principle be observed in experiments, so the $T=0$ ground state being superconducting is not of interest, or for that matter a $T_{c}$ which is exponentially small and therefore ‘impractical’ is also not of interest in the current work. We discuss both 2D and 3D cases, focusing mostly on the 3D case. We find that almost any theory based on the BCS-Migdal-Eliashberg paradigm, with some form of screened Coulomb coupling replacing the electron-phonon coupling in the BCS or Eliashberg theory, would uncritically predict absurdly high $T_{c}\sim 100$ K for $s$-wave superconductivity in all metals (including the alkali metals, which are well-described by the jellium model) arising from the unavoidable fact that the Fermi, plasmon, and Coulomb potential energy scales are all $10^{4}$ K. Therefore, we conclude, based on reduction ad absurdum, that the violation of the venerable Migdal theorem in this problem is sufficiently disruptive that no significance can be attached to numerous existing theoretical publications in the literature claiming plasmon-induced (or other similar Coulomb coupling-induced) practical SC. Using a careful analysis of the Eliashberg gap equations we find that the superconducting $T_{c}$ of the 3D (or 2D) electron gas can be reduced well below $\sim 1$ K depending on choices of frequency cut-off parameters that are introduced to satisfy Migdall’s theorem but are apriori unknown. The only believable result is the one discovered sixty years ago [Kohn1965] by Kohn and Luttinger predicting non-$s$-wave SC arising from Friedel oscillations with exponentially (and unobservably) low $T_{c}$. We provide several theoretical approaches using both BCS and Eliashberg theories and different screening models to make our point.
@@ I Introduction
Superconductivity is a ubiquitous phenomenon which has remained central to physics for more than 100 years ever since its discovery in 1911 [Onnes1911]. There are currently roughly $\sim 15{,}000$ known superconductors with $T_{c}$ ranging from $\sim 10$ mK to $\sim 250$ K (under high pressure). The search for new superconducting materials with practical use is an active research area in materials science with many promises of quantum computers and large language model based artificial intelligence revolutionizing such efforts often hyped in the media. On a fundamental level, superconductivity, being a quintessential example of spontaneous symmetry breaking (namely, the $U(1)$ symmetry), continues to attract great attention from theorists and experimentalists alike with the ongoing recent focus being room temperature superconductors, topological superconductors, and exotic superconductors with unusual order parameter symmetries. It may not be an exaggeration to say that superconductivity may very well be the most actively studied single topic in all of physics over the last 100 or so years.
@@
In this context, one issue that has fascinated physicists a great deal over the years is the superconducting mechanism—what leads to the pairing glue causing the superconducting instability? The minimal physical picture for superconductivity, and this is the relevant physics for the current work too, is that electrons pair up (“Cooper pairs” [Cooper1956]) because of an attractive interaction coupling them at the Fermi level causing the exchange of a virtual boson which acts as the glue to create the Cooper pairs. The glued Cooper pairs then condense into a zero resistance collective ground state producing the breaking of $U(1)$ symmetry since the Cooper pairs form the condensate effectively breaking the gauge invariance. The key physics is, however, the pairing arising from an effective attractive interaction between opposite spin electrons at the Fermi surface. Thus, a mechanism is needed to cause an effective attractive interaction between electrons near the Fermi surface. For most superconductors (if not all) this pairing arises from the electron-phonon interaction, which leads to a phonon-mediated effective attraction between the electron pairs with opposite spins near the Fermi surface. Since the electron spins are opposite, the spin part of the paired wavefunction is antisymmetric, making the spatial symmetry to be symmetric, and hence a singlet $s$-wave superconductivity. All superconducting metals are such spin-singlet spatially symmetric $s$-wave superconductors, with phonons being the bosonic glue between the electron pairs. In fact, one could probably claim that there is no known superconductor which has been compellingly shown to have a pairing glue other than electron-phonon coupling although there certainly are candidate materials where the situation remains an open question such as high-$T_{c}$ cuprate superconductors, heavy fermion superconductors, iron based pnictide superconductors, and perhaps various superconductors arising in multilayer graphene-type systems.
@@
The purpose of the current work, which is partly a perspective, is to investigate critically the extent to which the electron-electron interaction by itself could lead to any ‘practical’ $s$-wave superconductivity without the presence of any electron-phonon interaction (or other bosonic coupling outside the scope of just the electron system itself) in the Hamiltonian. At first, the question seems absurd since the basic electron-electron interaction, i.e., Coulomb coupling, is by definition repulsive, seemingly ruling out prima facie any attractive glue for pairing, let alone the formation of a superconducting condensate of the pairs. It turns out, however, that the situation is subtle since electrons screen each other, and the actual interaction between electrons is not the bare Coulomb interaction, but the many-body screened interaction, whose structure in general is extremely complicated. A seminal paper [Kohn1965] by Kohn and Luttinger (KL SC) pointed out a long time ago that the ground state of an interacting electron liquid, at least for a weak short-range screened interaction, is superconducting, hinting at the possibility that interacting electron liquids may very well manifest superconductivity arising just from the Coulomb coupling. But the KL SC is neither the conventional $s$-wave SC nor ‘practical’ by any stretch of imagination, as the estimated $T_{c}$ is exponentially low for real metals ($T_{c}\sim 10^{-10}$ to $10^{-100}$ K), and occurs only for higher (i.e. not $s=0$) orbital angular momenta. Thus, KL SC is only of academic interest, establishing a matter of principle that indeed repulsive Coulomb coupling under some circumstances can lead to a superconducting ground state arising from the existence of a sharp Fermi surface, albeit a very fragile unconventional one of no practical interest. KL SC arises from the Friedel oscillations in screening associated with the existence of the Fermi surface which leads to $2k_{F}$-periodic oscillatory behavior in the statically screened Coulomb interaction, and as such can only occur not for the spherically symmetric $s$-wave SC, but for higher angular momentum orbital symmetry such $p$, $d$ wave SC. There has been renewed interest in KL SC in the context of lattice systems such as the Hubbard model [Raghu2010] and 2D moiré materials [Chou2025] where the existence of van Hove singularities may enhance the KL SC $T_{c}$, but it is still low and the SC is always in higher orbital angular momentum channel. We mention as an aside, however, an important point not emphasized often in the context of KL SC in 2D moiré materials: the Kohn-Luttinger mechanism is likely not operational in a flat band (with an effective mass of infinity) since there is no sharp Fermi surface and no Friedel oscillations in a strict flat band.
@@
So, the KL SC, while establishing a matter of principle in the existence of SC via a strictly non-phonon mechanism, does not address the question we pose in this work: Can there be conventional $s$-wave metallic superconductivity which is also practical (i.e. experimentally observable with a reasonable $T_{c}$)? Much work has been done since the 1980s in the possibility of the so-called ‘plasmon-mediated’ metallic superconductivity in electron liquids, with claims of high $T_{c}$ $s$-wave metallic superconductivity arising from only electron-electron interactions. The basic physical idea is deceptively simple and appealing: Plasmons are quantized bosonic collective modes of the well-known ‘plasma wave’ charge density oscillations arising from the long-range Coulomb coupling, and the electron-plasmon interaction may act similarly to the electron-phonon interaction providing the pairing glue, leading to metallic superconductivity. If so, the putative $T_{c}$ of such plasmon-induced superconductivity could be very high since the energy scales of all electronic excitations in metals are extremely high, $10^{4}$ K in contrast to phonons whose characteristic energy scale is $\sim 10^{2}$ K. So, by just changing the bosonic glue from phonons to plasmons, it seems plausible that the $T_{c}$ could be raised by >2>2 orders of magnitude! In fact, the cuprate SCs with $T_{c}\sim 100$ K are thus heuristically consistent with an effective electron-electron induced SC since their $T_{c}$ is crudely a factor of 100 larger than the corresponding $T_{c}\sim 1$$10$ K in conventional metals such as Al and Pb which are known to manifest phonon-mediated SC for sure. This has motivated people to propose the plasmon mechanism for the cuprate superconductivity although this is definitely a rather small minority view by no means accepted in the larger high-$T_{c}$ community [Ruvalds1987, Ishii1993, Bill2003, Kresin1988]. The crucial problem in such suggestions for plasmon mediated superconductivity in electron liquids is the stark absence of generic high-temperature ($T_{c}>100$ K) metallic superconductivity in normal metals which all have plasmons with energy scales $10^{4}$ K and Fermi energies also $10^{4}$ K. If plasmons and electron-electron interactions could generically lead to metallic superconductivity, Na and K and Li should all be room temperature superconductors, but factually, they are not superconductors at all under ambient conditions in spite of being ideal jellium electron liquids with electronic energy scales $\sim 5\times 10^{4}$ K. But, all known (>50>50 different elements, some under high pressure) metallic superconductors with plasmon energies $10^{4}$ K, are known to be phonon-mediated $s$-wave SCs with $T_{c}$ (mostly) $\sim 1$$10$ K which is in line with the typical phononic energy scale $\sim 10^{2}$ K. In fact, a major problem with the plasmon-mediated (or other similar) electron interaction mediated SC proposal is: Why is the $T_{c}$ so low in all such SCs (even in high-$T_{c}$ cuprates, $T_{c}\sim 100$ K, orders of magnitude below the electronic kinetic energy and interaction energy scales)?
@@
We show in this work, using plausible BCS theory based arguments, that the generically expected BCS $T_{c}$ for the plasmon mediated metallic SC is $\sim 100$$1000$ K, an absurdly high number when talking about conventional metallic superconductors which have been experimentally studied extensively for almost 120 years. Obviously, this ‘plausible’ straightforward BCS theory, with the plasmon modes replacing the phonon modes and the electron-plasmon coupling replacing the electron-phonon coupling, does not work as it predicts manifestly incorrect $T_{c}$ which is incompatible with reality. Alkali metals are not $\sim 100$ K superconductors, and the superconducting metals get their superconductivity from electron-phonon interactions, not the electron-electron interactions. Therefore, the question arises what is wrong with the BCS theory applied to plasmons compared with that for phonons where generally it has had predictive success in producing $T_{c}$ and superconducting gaps in agreement with experiments in numerous situations [Allen1983, Allen1999]. We discuss this point from several critical perspectives using the Eliashberg theory, emphasizing that it is possible to predict essentially any $T_{c}$ based on the BCS-Eliashberg theory for plasmon-induced superconductivity since the theory is uncontrolled because of the crucial inapplicability of Migdal theorem for the electron-plasmon interaction [Migdal1958]. The successful BCS-Eliashberg theory for superconductivity, enabling a reasonable quantitative $T_{c}$ prediction for metallic superconductors, is grounded on the Migdal theorem which applies to the electron-phonon interaction problem in metals by virtue of the Fermi energy being much larger than the typical phonon energy (the so-called Debye temperature), but not for the electron-electron interaction induced plasmonic superconductivity.
@@
Now we immediately face a serious conundrum in discussing superconductivity induced by plasmons. Since the plasmons arise intrinsically as collective modes of the interacting electron system, unlike the phonons which arise from the lattice vibrations, there cannot be, by definition, any Migdal theorem for the electron-plasmon interaction since there is no separation between electron and plasmon energy scales. Indeed, the plasmon energy in metals ($\sim 10$ eV) is comparable to the Fermi energy $E_{F}$, and is also similar to the typical average metallic inter-electron Coulomb potential energy. Thus, there is no energy scale separation (or velocity scale separation) enabling the neglect of vertex corrections. Hence, the use of BCS-Eliashberg theory for the calculation of $T_{c}$ in the plasmon induced metallic superconductivity is fundamentally flawed as one must include vertex corrections to all orders in the theory, which is obviously impossible. We show that it is easy to obtain $T_{c}\sim 100$$1000$ K in metals arising from electron-plasmon interaction simply by following the standard BCS-Eliashberg-Migdal theory (with plasmons replacing phonons), but these results are unreliable since the neglect of vertex corrections in the theory is unjustified and uncontrolled. Similarly, the estimate of $\mu^{*}$ in the plasmon case is nontrivial and ambiguous since one is starting entirely with a repulsive interaction. We elaborate on these subtle but important theoretical issues in the rest of this paper.
The rest of this paper is organized as follows. In Sec. II, we provide the calculated dynamically screened Coulomb interaction in 3D and 2D jellium metals for several many-body approximations to show explicitly that indeed the effective screened interaction is attractive in large regimes of the energy-momentum space, making a discussion of possible superconductivity arising strictly from electron-electron interactions meaningful. In Sec. III, we use the effective electron-plasmon interaction from Sec. II to obtain the SC $T_{c}$ induced by plasmons in the standard BCS theory, finding $T_{c}\sim 100$ K in typical metals (which is in stark disagreement with experiments). In Sec. IV, we develop a full Eliashberg-Migdal theory, and point out precisely why the theory is unreliable in predicting $T_{c}$ by discussing the subtle frequency-dependent retardation effects which are crucial in the Eliashberg theory. We conclude in Sec. V, summarizing our findings and emphasizing that no compelling theoretical case can be made that electron-plasmon interaction can or should lead to conventional and practical superconductivity in normal metals, although uncontrolled approximations do typically lead to $T_{c}\sim 100$ K for plasmon-induced metallic superconductivity. Appendix A provides some of the technical details on the screened interaction.